7. You can purchase a 10 year,4.250% coupon note that pays interest quarterly and has a quoted yield to maturity of 3.755%. However, you prefer to invest in semiannual coupon bonds.

(a) What quoted yield to maturity on a semiannual coupon bond would make you indifferent between that bond and the quarterly coupon bond described above?

(b) There is a special name for the yield to maturity on a semiannual coupon bond. What is it?
Answer #1

1.
(1+semiannual quoted rate/2)^2=(1+quarterly quoted rate/4)^4
Hence,
(1+r/2)^2=(1+3.755%/4)^4
=>r=((1+3.755%/4)^(4/2)-1)*2
=>r=3.7726%

2.
Bond Equivalent Yield